How do I solve |3x+1|=|3x-1|?
Answer by josgarithmetic(39618)

Very briefly, either they are opposite in sign or the same in sign.

3x+1=-(3x-1)
3x+1=-3x+1
6x=0
x=0----the only answer.

OR

3x+1=3x-1
1=-1, false.
No Solution.